

********************************************************************
UARS WEEKLY STATUS REPORT - 08 JUN 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM has performed its daily solar scans for this past week
during periods of predicted UARS telemetry.
***********
ACRIM
***********
ACRIM II exhibited some out of range indicators at the FOT in early May.
The FOT powered down the instrument on 5/5/01 in response to their internal
rules and we have not been able to power the instrument back up. We have
confirmed that good data was being acquired at the time of the power-down.
An anomaly investigation has been convened by the UARS Project to explore
the problem and identify possible solutions.
Re-processing ACRIM II data has been completed and the Version 2 results
compiled through 5/5/01. These are available on the ACRIM web site
(http://acrim.com) in tabular and graphical formats. They will be submitted
to the GSFC and LaRC DAAC's as well as the NOAA NGCD shortly.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating region. Science data products are of excellent
quality.
A number of papers were presented at the Spring AGU Meeting of May/June
2001 in Boston that included analyses of various aspects of the HALOE
dataset. In particular, there were Special Sessions on Polar Mesospheric
Clouds and on Middle Atmospheric Odd Nitrogen. A number of the
presentations in those sessions were focused on the HALOE data. One new
HALOE finding, reported by Dr. Mark Hervig of the U. of Wyoming, involved
the interpretation of spectral observations of cloud tops just above the
cold tropical tropopause. Those clouds frequently have a clear signature
of nitric acid trihydrate (NAT).
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in a UARS Battery Status Review
Meeting on 4 June 2001 and the UARS Monthly Status
Review Meeting on 6 June 2001. Daily processing of
instrument Daily Activity Plans continues with no problems.
***********
CDHF
***********
- DAAC DATA:
- Final approval has been received from the CLAES group to load the
Version 9 into the DAAC Transfer Tables for data created in the
month of June, 1999.
- Data created in the month of March, 2001 for HALOE, NMC, and UKMO
has been loaded into the DAAC Transfer Tables. DAAC has been notified.
- UARS CDHF REHOST:
- Software to "quarantee puts to Silo" is operational. We are running
test cases to validate software.
- SUSIM 7.2 software was reinstalled on the rehost system (25-MAY-2001)
to allow assess to the new version of UOAS (CDS_VIEW_VEC_CHECK).
***********
FOT
***********
REPORTING PERIOD
This report covers 26 May 2001 (Orbit 53056, GMT Day 146) through 01 June 2001
(Orbit 53160, GMT Day 152).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 146/2356 53070 Power OFF
PEM 148/2259 53100 Power OFF
SOLSTICE 148/2357 53100 Day Only
ACRIM 148/2357 53100 Power OFF
SUSIM 148/2357 53100 Power OFF
SOLSTICE 150/1018 53122 Power OFF
A Forward-to-Reverse Yaw Around maneuver (number 99) was successfully executed
in Orbit 53125 beginning 150/15:16:50 and completing at 150/15:54:00 on 30 May
2001. The SOLSTICE instrument was in day only and all other instruments were
in their off or standby mode (including HRDI and PEM). Due to the SA position
at 269 degrees, the effective night length of the spacecraft after the maneuver
was approximately 54.6 minutes. Post maneuver analysis of data indicate that
during the Yaw Around, the end-of-night load bus voltage was 24.00 volts and
the average depths of discharge (DODs) on Batteries 2 and 3 were approximately
22.0 and 23.5 percent respectively. The next (reverse-to-forward) YAW is
scheduled for 28 June 2001 (DOY 179).
The beta angle ranged between the angles of 16.0 degrees to 6.9 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 34 degrees on 14 June 2001 (DOY 165).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 to 1.3 degrees C. Additional battery performance data for each
battery on 25 May 2001 and 01 June 2001 are provided in Appendix D.
The clock error ranged from -13.3 to 17.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 17
¥ SUSIM = 02
¥ PEM = 01
¥ ACRIM = 01
¥ HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/137 show the observatory data loss to be 301 hours, 38
minutes, 17.048 seconds. This shows a loss of 22 minutes and 6.760 seconds
since the last report period. This is a 0.39 percent data loss, which equals a
99.61 percent data capture for the mission.
The increase of 22 minutes and 6.760 seconds occurred over four (4) days:
· 37.632 seconds on Day 147 due to a Data Hit.
· 13.152 seconds on Day 149 due to a Data Hit.
· 21 minutes and 01.880 seconds on Day 150 due to a corrupt stat table in MA
Subs. Controller.
· 14.096 seconds on Day 151 due to a Data Hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to fine-tune the TAM alignment by loading new alignment
matrices. A new alignment is required for the reverse flight direction. This
is a slow process and it will take many alignments to lock in on the most
accurate solution.
An instrument anomaly was experienced on 05 May 2001. The instrument involved
was the ACRIM II instrument. All ACRIM telemetry for currents and voltages all
dropped to zero values suddenly. The ACRIM team in coordination with the
UARSFOT is conducting analysis and command tests to determine the cause of the
anomaly. Preliminary analysis shows that the anomaly was caused due to failed
circuitry in the ACRIM electronics box.
MEETINGS
The FOT attended a UARS TAC RS Meeting on Wednesday, May 30, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 28 June 2001 (DOY 179). This maneuver
will be a Reverse to Forward Flight maneuver (Southern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 01 June 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
53056 146/0037 01-336 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
53070 146/2356 01-337 HALOE Power OFF
53070 146/2356 01-338 OBC/CLK Clk Adj 20.6 to 0.00 msec/day
53074 147/0505 01-339 NBTR PBK / RTS-58
53074 147/0505 01-340 OBC/CLK Clk Adj 0.00 to -20.6 msec/day
53095 148/1521 01-341 OBC/CLK Clk Adj -20.6 to 0.00 msec/day
53100 148/2259 01-342 PEM DAY ONLY TO OFF
53100 148/2357 01-343 SOLSTICE Power OFF
53100 148/2357 01-344 ACRIM Power OFF
53100 148/2357 01-345 SUSIM Power OFF
53102 149/0322 01-346 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
53111 149/2030 01-347 NBTR NOOP
53116 150/0139 01-348 OBC/CLK Clk Adj 20.6 to 0.00 msec/day
53120 150/0550 01-349 OBC/CLK Clk Adj 0.00 to -20.6 msec/day
53122 150/1018 01-350 SOLSTICE DAY ONLY TO POWER OFF
53124 150/1420 01-351 MACS TAM TO MON. MODE
53124 150/1420 01-352 MACS YAW TO REV
53125 150/1548 01-352 MACS YAW TO REV
53125 150/1548 01-353 MACS TAM TO UPDATE
53126 150/1616 01-352 MACS YAW TO REV
53139 151/1436 01-354 MACS TAM1
53145 151/2316 01-355 OBC Clk Adj fr -20.6 to 0.00 msec/day
53146 152/0056 01-356 OBC Clk Adj fr 0.00 to 20.6 msec/day
______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 01 June 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 01 June 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
53056 146/0050 Clk Adj 0.00 to 20.6 msec/day
53070 147/0012 Clk Adj 20.6 to 0.00 msec/day
53074 147/0523 Clk Adj 0.00 to -20.6 msec/day
53095 148/1521 Clk Adj -20.6 to 0.00 msec/day
53102 149/0322 Clk Adj 0.00 to 20.6 msec/day
53116 150/0148 Clk Adj 20.6 to 0.00 msec/day
53120 150/0556 Clk Adj 0.00 to -20.6 msec/day
53145 151/2318 Clk Adj -20.6 to 0.00 msec/day
53146 152/0058 Clk Adj 0.00 to 20.6 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 01 June 2001
25 May 2001 - GMT Day 145 Beta = 21.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-106.4 +5.79/+4.51 +29.6/-14.4 24.6 1.040 21.0
3 +22.4/-56.0 +4.51/+3.23 +28.8/-14.4 24.5 1.030 20.9
1 June 2001 - GMT Day 152 Beta = 6.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/+0.0 12.0 0.000 0.0
2 +22.4/-95.2 +5.47/+4.51 +28.8/-13.2 24.5 1.030 20.9
3 +22.4/-78.4 +4.51/+3.23 +28.0/-13.6 24.5 1.020 20.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
___________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 01 June 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-031 FOT 53127 150/1909 23913 Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-025 FOT see list May 01 14170 Generic Late Acquisition
01-031 FOT 53127 150/1909 23913 Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-022 ACRIM 52709 122/2140 N/A Red Low voltage alert
01-024 ACRIM 52753 125/2045 N/A ACRIM Alerts - Power Off
SEVNTFW01/22